OpenStudy (anonymous):

Given the table below, determine if the data represents a linear or an exponential function and find a possible formula for the function. x 0 1 2 3 4 f (x) 12.5 13.75 15.125 16.638 18.301

OpenStudy (anonymous):

If a function is linear, the y values will increase/decrease by a number that is added to the previous factor. If a function is exponential, the y values will increase/decrease by a number that the previous factor is multiplied by. Linear: x, x+2, x+4, x+6 Exponential: x, 2x, 4x, 8x

OpenStudy (anonymous):

What do i do next?

OpenStudy (anonymous):

If you know what type of function you have, you write a general formula. What type of function is this?

OpenStudy (anonymous):

it's linear

OpenStudy (anonymous):

noo exponential

OpenStudy (anonymous):

good, its exponential because you multiply each y value by a factor of 1.1 to get to the next y value. Now you write the general formula for the sequence.

OpenStudy (anonymous):

The general formula for a geometric (exponential) sequence is the number that you are multiplying one y value by to get to the next one to the power of the corresponding x value plus some number. So what you have so far is y=1.1^(x+?) You can solve for your unknown by plugging in an (x,y) Its easiest to use 0 as your x, so we'll plug in the coordinate (0,12.5) 12.5=1.1^(0+?) Now to find the value of the unknown, you write a log log(12.5) = (0+?)log(1.1) divide both sides by log 1.1 0+? = approximately 26.5 You use this number for every equation, and the x and y values change, so if you wanted to find the y value for x=5, you would do: y=1.1^(5+26.5) y=1.1^(31.5) y= 20.131